What is the distance between 3kg and 5kg mass particle if the gravitational force on 3kg due to 5kg particle is 5.12x10^-6 N?


1
Expert's answer
2021-11-15T10:32:58-0500
"F=\dfrac{Gm_1m_2}{d^2},""d=\sqrt{\dfrac{Gm_1m_2}{F}},""d=\sqrt{\dfrac{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}\times3\ kg\times5\ kg}{5.12\times10^{-6}\ N}}=1.4\times10^{-2}\ m."